College Physics 1e LS W 12 12 (7 of 15) e Time Sp 00:02: 7 8910 At what depth under the surface of gasoline is the pressure the same as that at a depth h under the surface of water? Express your answer as a multiple of h. 1.38 I have tried 1.31 1.32 1.33 1.34 1.35 1.36 1.37 1.38 1.39 and 4.0 as all answers and none are correct... not sure what I'm doing wrong
media%2F094%2F094f7062-8bb0-4f3d-aa2d-67
Answer #1


In water,pressure at a depth is P = Pa+(rho*g*h)

Pa is the atmospheric pressure = 1.013*10^5 Pa

rho is the density of water

then

P = Pa+(1000*9.8*h)

In gasoline ,density is rho_g = 719.7 kg/m^3


Pg = Pa+(719.7*9.8*h)


then given that


P = Pg

Pa+(1000*9.8*h) = Pa+(719.7*9.8*h')


h' = (1000*h)/719.7 = 1.39*h